OpenStudy (anonymous):

Prove, in complete sentences, whether the left expression is equal to the right expression and discuss which property applies. 2(x − 3) <-> 2x − 6 I know that they're equal (or I think anyway), but I'm not sure what it means by what properite applies.

OpenStudy (ivettef365):

a(b+c) = ab+ac

OpenStudy (anonymous):

3/(2x) + x/(2x - 6) = 3/(2x) + x[2(x - 3)] = [3(x - 3) + x(x)]/[2x(x - 3)] = (3x - 9 + x^2)/[2x(x - 3)] = (x^2 + 3x - 9)/[2x(x - 3)

OpenStudy (ivettef365):

that is distributive

OpenStudy (anonymous):

yea there equal

OpenStudy (anonymous):

yea like ivette said, the left side you distribute, times 2 to everything.

OpenStudy (anonymous):

hen they say which properties apply your suppose to put distribuive property

OpenStudy (anonymous):

when*

OpenStudy (anonymous):

distributive property*

OpenStudy (anonymous):

you understand??

OpenStudy (anonymous):

Do you want a REALLY long explanation?

OpenStudy (anonymous):

this will help http://www.geteasysolution.com/2(x-3)=2x-6

OpenStudy (anonymous):

Because: Solve for x: 2 (x-3) = 0 Divide both sides by a constant to simplify the equation. Divide both sides of 2 (x-3) = 0 by 2: (2 (x-3))/2 = 0/2 Any nonzero number divided by itself is one. 2/2 = 1: x-3 = 0/2 Any number times zero is zero. 0/2 = 0: x-3 = 0 Isolate terms with x to the left hand side. Add 3 to both sides: x+(3-3) = 3 Look for two terms that sum to zero. 3-3 = 0: Answer: | | x = 3 ^ That's the first equation, and this is the second equation: Solve for x: 2 (x-3) = 0 Divide both sides by a constant to simplify the equation. Divide both sides of 2 (x-3) = 0 by 2: (2 (x-3))/2 = 0/2 Any nonzero number divided by itself is one. 2/2 = 1: x-3 = 0/2 Any number times zero is zero. 0/2 = 0: x-3 = 0 Isolate terms with x to the left hand side. Add 3 to both sides: x+(3-3) = 3 Look for two terms that sum to zero. 3-3 = 0: Answer: | | x = 3 Just explain both are processed the same way, with the same solution.

OpenStudy (anonymous):

omg!

OpenStudy (anonymous):

Ok, thanks everyone for all of your great answers. I think I have it now. How do you think this sounds? Prove, in complete sentences, whether the left expression is equal to the right expression and discuss which property applies. 2(x − 3) <-> 2x − 6 The left expression is equal to the right expression, because 2(x - 3) = 2x - 6. Therefore, 2(x - 3) <-> 2x - 6. The distributive property applies here.

OpenStudy (anonymous):

perfect

OpenStudy (anonymous):

You got it :)

OpenStudy (anonymous):

Okay, thanks!

OpenStudy (anonymous):

Oh, one more question, 6(2r − 5) + 2r simplified would be 14 r − 30, right?

OpenStudy (anonymous):

Starting Problem: 6(2r - 5) + 2r= ? Step 1: Multiply multiply the 6, 6(2r - 5) + 2r = (6 x 2r - 6 x 5) + 2r = (12r - 30) + 2r Step 2: Combine like terms = 12r - 30 + 2r = 12r +2r - 30 = 14r - 30 Solution: 14r - 30

OpenStudy (anonymous):

does that help?

OpenStudy (anonymous):

Yes! Thank you! :)

OpenStudy (anonymous):

great